OpenStudy (anonymous):

. Find the sum of the geometric sequence. (1 point) 1, 1/2, 1/4, 1/8, 1/16.

OpenStudy (anonymous):

@nincompoop any ideas?

Nnesha (nnesha):

ther is 2 formulas to find sum of geothermic sequence : sum of finite geothermic seq. and sum of infinite geothermic seq.

Nnesha (nnesha):

The formula for the sum of finite geometric seq. \[\huge\rm s_n = a_1 (\frac{ 1 -r^n }{ 1-r })\] if r is less than one then use this formula for infinite geo seq. \[\huge\rm s = \frac{ a_1 }{ 1-r }\] ^^^ if \[\left| r \right| < 1\] r = ratio

OpenStudy (anonymous):

Alright @Nnesha how would I use that

Nnesha (nnesha):

first find ratio

Nnesha (nnesha):

do you know how to find r ??

OpenStudy (anonymous):

No

Nnesha (nnesha):

r is the common ratio of the sequence you can find r by dividing next term with the previous one \[\huge\rm r = \frac{ a_2 }{ a_1 } = \frac{ a_4}{ a_3}\] so on

Nnesha (nnesha):

\[a_1\] first term \[a_2\] 2nd term

OpenStudy (amistre64):

a geometric sequence is defined by some common value is multiplied each time to get the next value. so: 1*r = 1/2 should tell us the value of r

Nnesha (nnesha):

yep that's right to find ratio you cane either multiply or divide both ways you will get same answer for example 3 ,9 ,27 3 r = 9 to solve for r you have to DIVIDE both side by 9 so r =9/3 ~~~~ which is same as r =a_2/a_1